OpenStudy (anonymous):

Word problem Help: Two trucks can carry loads in a ratio of 5 to 2. The smaller truck has a capacity 3 tons less then that of the larger truck. What is the capacity of the larger truck?

OpenStudy (anonymous):

X-3/x =2/5 , solve for x , you get x= 5

OpenStudy (anonymous):

Thank you very much